A particular stock has a dividend yield of 0.6 percent. Last year, the stock price fell from $71 to $52. What was the return for the year? (A negative value should be indicated by a minus sign. Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)
Answer #1

Capital gains yield=(End value-Beginning value)/Beginning value

=(52-71)/71=(26.76%)(Approx)(Negative).

Hence return for the year=Dividend yield+Capital gains yield

=(0.6-26.76)

=(26.16%)(Approx)(Negative).